# Jason Gaitonde

### v1

@@ 1.1 Related Work
To our knowledge, a general, tight comparison result has not yet been established. Very recent work of Chlebicka, Łatuszyński, and Miasodejow [CŁM24] establishes a quantitative comparison of spectral gaps: if the Glauber dynamics has spectral gap of order $1/n^{\beta}$, then a full round of any systematic scan has spectral gap at least $\Omega(1/n^{2\beta+2})$. Their work also views the site transitions as orthogonal projections and employs linear-algebraic machinery of Badea, Grivaux, and Müller [BGM11] on convergence of alternating projections. Our approach is significantly more elementary, direct, and yields a tight spectral gap lower bound of $\Omega(1/n^{\beta+1})$ in their setting. We also provide a qualitative converse that establishes a polynomial relationship in the reverse direction, answering Open Question 3.7 of their work. The results of He, de Sa, Mitliagkas, and Ré [HSMR16] provides a comparison result between the lazy systematic scan and Glauber dynamics that is qualitatively and quantitatively weaker, with a polynomial dependence on the minimal holding probability which may be exponentially small even in rapid mixing spin glasses.
The special case where $n=2$ has been significantly better studied since the scan remains reversible and can be viewed as alternating projections. As demonstrated by Roberts and Rosenthal [RR15], several phenomena that arise are particular to this case and cannot extend to larger $n$, so the techniques and results do not appear to have much bearing for the general problem we consider. At the linear-algebraic level, a version of the matrix AM-GM inequality is known to hold [BK00] when $n=2$; this implies that the spectral gap of the scan is at least as large as that of Glauber. This is intuitively clear for the Gibbs sampler, since there is no benefit from applying a site transition twice in a row. Several more precise results for this special case have been established in the literature, see e.g. [DKSC10, And16, GKZ18, QJ22] for more discussion on this setting.
@@ 2.1 Markov Chains, Gibbs Samplers and Linear Algebra
$\mathsf{P}_{i}f(\bm{x})=\mathbb{E}_{X\sim\pi}[f(X)|X_{-i}=\bm{x}_{-i}].$ (1)
It is well-known that $\mathsf{P}$ is self-adjoint and satisfies $\mathsf{P}_{i}^{2}=\mathsf{P}_{i}$. The former is particularly easy to see in the case each $\mathcal{X}_{i}$ is finite, in which case one can take $\mu_{i}$ to be uniform and so if $\bm{y}=(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{n})$,
@@
We may now define the relevant operators that will be the subject of this work. For the rest of the paper, we assume that each $\mathsf{P}_{i}$ is given by 1.
@@ 2.2 Mixing Times
For reversible Markov chains, the mixing times and operator norm are closely related to the eigenvalues, which all exist and are real by virtue of the spectral theorem. One can provide an analogous lower bound on the mixing time for non-reversible chains in terms of the spectrum, but this is analytically quite difficult since they lie in the unit complex disc. However, we have the recent lower bound of Chatterjee [Cha23] that will prove useful in providing a full comparison for our setting:
Theorem 2.2 ([Cha23], Theorem 1.4).
@@ 3.1 From Glauber to All Scans
Note that Theorem 3.1 holds under any other (non-identity) ordering in 3 of first appearances of the set of indices in $\mathsf{P}_{i_{L}}\ldots\mathsf{P}_{i_{1}}$ by symmetry; we choose to write it for the identity ordering simply to avoid additional notation. The following is immediate:
@@
Let $\gamma_{\mathsf{GD}}>0$ denote the spectral gap of any Glauber dynamics and $\gamma_{\mathsf{SS}(\sigma)}$ the spectral gap of the systematic scan with permutation $\sigma$. Corollary 3.1 instantiated with the Gibbs transition matrices implies
@@
These increments are independent, and so
@@
Fix any $f\in\mathcal{H}$ satisfying $\|f\|=1$ and define the sequence $0:=\gamma_{0}\leq\gamma_{1}\leq\ldots\leq\gamma_{L}$ via
@@
$\leq\sum_{\ell=1}^{k_{j}-1}\left\|\mathsf{Q}_{{\ell-1}}f-\mathsf{Q}_{{\ell}}f)\right\|+\sum_{\ell=1}^{k_{j}}\left\|\mathsf{Q}_{{\ell}}f-\mathsf{Q}_{{\ell-1}}f\right\|$
@@
Taking squares, applying Cauchy-Schwarz, and telescoping via 4, we find that
@@
By assumption, the left hand side is at most $(1-\delta)^{2}\leq 1-\delta$ so rearranging yields
@@ 3.2 From Scans to Glauber
We now show how a variation of the argument of Theorem 3.1 can be used to show a converse: if all, or even any, systematic scans have a spectral gap, then the Glauber dynamics has a spectral gap that is polynomially related. This provides a simple resolution to Open Problem 3.7 of Chlebicka, Łatuszyński, and Miasodejow [CŁM24]. However, unlike Theorem 1.1, we expect that our quantitative dependence is quantitatively suboptimal.
@@
$\mathsf{Q}_{{L-1}}f=\mathsf{Q}_{{j_{n-1}}}f+\sum_{\ell=j_{n-1}}^{L-2}(\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f),$
so a similar set of manipulations with the triangle inequality, Cauchy-Schwarz, and the fact projections can only decrease distance (noting $i_{j_{n-1}}=n-1$ by definition), and 7 yields
$\|\mathsf{Q}_{{L-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|$ $=\left\|\mathsf{Q}_{{j_{n-1}}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f+\sum_{\ell=j_{n-1}}^{L-2}(\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f)\right\|$
$\leq\|\mathsf{Q}_{{j_{n-1}}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|+\sum_{\ell=j_{n-1}}^{L-2}\|\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f\|$
$=\|\mathsf{P}_{n-1}\mathsf{Q}_{{j_{n-1}-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|+\sqrt{j_{n}-j_{n-1}-1}\sqrt{\sum_{\ell=j_{n-1}}^{L-2}\|\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f\|^{2}}$
$\leq\|\mathsf{Q}_{{j_{n-1}-1}}f-\mathsf{P}_{n-2}\ldots\mathsf{P}_{1}f\|+\sqrt{j_{n}-j_{n-1}-1}\sqrt{\sum_{\ell=j_{n-1}}^{L-2}(\lambda_{\ell+1}-\lambda_{\ell})}.$
@@
$\|\mathsf{Q}_{{L-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|$ $\leq\sum_{k=1}^{n-1}\sqrt{j_{k+1}-j_{k}-1}\sqrt{\sum_{\ell=j_{k}}^{j_{k+1}-2}(\lambda_{\ell+1}-\lambda_{\ell})}$
$\leq\sqrt{\sum_{k=1}^{n-1}(j_{k+1}-j_{k}-1)}\sqrt{\sum_{\ell\in S}(\lambda_{\ell+1}-\lambda_{\ell})}$
@@
where $S$ is a suitable subset of $[L]$; note that we use the fact that the inner summands for the increments are disjoint across different $k$ by construction. Putting this back with 9, we have
@@
$t^{\mathsf{GD}}_{\mathsf{mix}}(\varepsilon)\leq\max_{\sigma}\widetilde{O}\left(n^{4}(t_{\mathsf{mix}}^{\mathsf{SS}(\sigma)}(1/10))^{2}\right)\cdot\log\left(\frac{1}{\varepsilon\pi_{\min}}\right)$
$t^{\mathsf{GD}}_{\mathsf{mix}}(\varepsilon)\leq\min_{\sigma}\widetilde{O}\left(n^{6}(t_{\mathsf{mix}}^{\mathsf{SS}(\sigma)}(1/10))^{2}\right)\cdot\log\left(\frac{1}{\varepsilon\pi_{\min}}\right).$
@@ 4 Optimality of Results
Let $\pi$ be the uniform distribution on the set of spin configurations $\bm{x}\in\{0,1\}^{n}$ such that $\sum_{i=1}^{n}x_{i}\leq 1$. This is the so-called hardcore model of independent sets on the complete graph $K_{n}$ on $n$ nodes with fugacity $\lambda=1$. In particular, the set of configurations with positive (and uniform) probability is $\{\bm{0},\bm{e}_{1},\ldots,\bm{e}_{n}\}$, where $\bm{e}_{j}$ is the $j$th standard basis vector in $\mathbb{R}^{n}$. The associated Gibbs sampler is easily seen to act on configurations as follows when re-randomizing a spin $i$: if $X^{t}\in\{\bm{0},\bm{e}_{i}\}$, $X^{t+1}$ is uniform in $\{\bm{0},\bm{e}_{i}\}$, while otherwise $X^{t+1}=X^{t}$.
@@
If 12 holds, this implies the total variation between $c^{\prime}n^{2}$ full systematic scans and $\pi$ exceeds $1/2$ since $\pi(\{\bm{e}_{i}:i\in A\}\cup\{\bm{0}\})\leq 1/4$ while the dynamics has probability at least $3/4$ of being in this set.
To prove 12, observe that by independence of the geometric random variables in 11,
@@
$\Pr\left(\max_{s\in[s_{-},s_{+}]}\left|\sum_{i=1}^{s}Y_{i}-2s\right|\geq\lambda^{*}\right)\geq 1/10.$ (14)
Putting together 13 and 14 implies that
@@
Since $[2s_{-}-\lambda^{*},2s_{+}+\lambda^{*}]$ is a contiguous interval of length $O(\sqrt{c^{\prime}}n)$, the image $A$ of this set under the $\ (\mathrm{mod}\ n)$ map also has at most $n/4-1$ elements, and thus 15 demonstrates $A$ satisfies 12 so long as $c^{\prime}$ is taken to be a sufficiently small constant. By the discussion above, this verifies $t^{\mathsf{SS}}_{\mathrm{mix}}(1/4)\geq c^{\prime}n^{2}\geq cnt_{\mathrm{mix}}^{\mathsf{GD}}(1/4)$ so long as $c<c^{\prime}/C_{\mathsf{GD}}$.

### v2

@@ 1.3 Related Work
To our knowledge, our work is the first to both establish a tight comparison of the two samplers in one direction as well as a polynomial equivalence of the two samplers in general. The main result of Chlebicka, Łatuszyński, and Miasodejow [CŁM25] yields the following suboptimal guarantee: if the Glauber dynamics has spectral gap of order $\gamma$, then a full round of any systematic scan has spectral gap at least $\gamma^{2}/n^{2}$, which is quadratically worse than our optimal bound in both components. Consequently, their result cannot establish the (weaker) claim that the mixing time of systematic scans are at most $\mathsf{poly}(n)$ factors worse of Glauber dynamics in low-temperature systems with superpolynomial mixing times, nor ever prove a mixing time bound better than $\Theta(n^{5})$ steps for the systematic scan in any nontrivial system where $\gamma\lesssim 1/n$. At the technical level, their work employs machinery of Badea, Grivaux, and Müller [BGM11] on the convergence of alternating projections. Our approach is conceptually cleaner and more direct, yielding the optimal spectral gap. We also provide a qualitative converse that establishes a polynomial relationship in the reverse direction, answering Open Question 3.7 of their work using variations of this technique, emphasizing the applicability of our methods. By contrast, their results (as proven in earlier work of [RR97]) only establish an exponentially small spectral gap for the converse exploiting the trivial fact that expanding $n$ steps of Glauber includes a single systematic scan (out of all $n^{n}$ choices of sites) with an assumed spectral gap. We use our more flexible and direct framework for analyzing spectral gaps to establish the first polynomial relationship in this direction.
The earlier work of He, de Sa, Mitliagkas, and Ré [HSMR16] provides a litany of examples showing no asymptotically general ordering of scan dynamics can hold, and then provide a comparison result using conductances. However, this argument requires them to compare a lazy, augmented systematic scan with the Glauber dynamics, where the augmented state space is enlarged to incorporate the index of the site to be updated at the current state. Their result shows that the mixing time of Glauber can be bounded by the essentially the square of the mixing time of a lazy, augmented scan. On the other hand, their mixing time bound for the lazy, augmented scan only yields an upper bound in terms of Glauber that is quantitatively suboptimal in terms of additional $\mathsf{poly}(n,\tau_{\mathsf{mix}})$ factors and further scales inversely with the minimum holding probability of any transition, which can be exponentially small even in rapidly mixing models like high-temperature spin glasses.
The special case where $n=2$ can be viewed as alternating projections and admits considerably sharper guarantees. As demonstrated by Roberts and Rosenthal [RR15], several phenomena that arise are particular to this case and cannot extend to larger $n$, so the techniques and results do not appear to have much bearing for the general problem we consider. At the linear-algebraic level, a version of the matrix AM-GM inequality is known to hold [BK00] when $n=2$; this implies that the spectral gap of the scan is at least as large as that of Glauber. This is intuitively clear for the Gibbs sampler, since there is no benefit from applying a site transition twice in a row. Several more precise results for this special case have been established in the literature, see e.g. [DKSC10, And16, GKZ18, QJ22] for more discussion.
@@ 2.1 Markov Chains, Gibbs Samplers and Linear Algebra
$\mathsf{P}_{i}f(\bm{x})=\mathbb{E}_{X\sim\pi}[f(X)|X_{-i}=\bm{x}_{-i}],$ (5)
where we write $-S$ to denote the restriction to the coordinates in $[n]\setminus S$. It is well-known that $\mathsf{P}_{i}$ is self-adjoint and satisfies $\mathsf{P}_{i}^{2}=\mathsf{P}_{i}$. The former is particularly easy to see in the case each $\mathcal{X}_{i}$ is finite, in which case one can take $\mu_{i}$ to be uniform and so if $\bm{y}=(x_{1},\ldots,x_{i-1},y,x_{i+1},\ldots,x_{n})$,
@@
We may now define the relevant operators that will be the subject of this work. For the rest of the paper, we assume that each $\mathsf{P}_{i}$ is given by 5.
@@ 2.2 Mixing Times
For reversible Markov chains, the mixing times and operator norm are closely related to the eigenvalues, which all exist and are real by virtue of the spectral theorem. One can provide an analogous lower bound on the mixing time for non-reversible chains in terms of the spectrum, but this is analytically quite difficult since they lie in the unit complex disc. However, we have the recent lower bound of Chatterjee [Cha25] that will prove useful in providing a full comparison for our setting:
Theorem 2.2 ([Cha25], Theorem 1.4).
@@ 3.1 From Glauber to All Scans
Note that Theorem 3.1 holds under any other (non-identity) ordering in 7 of first appearances of the set of indices in $\mathsf{P}_{i_{L}}\ldots\mathsf{P}_{i_{1}}$ by symmetry; we choose to write it for the identity ordering simply to avoid additional notation. The following is immediate:
@@
Let $\gamma_{\mathsf{GD}}>0$ denote the spectral gap of the Glauber dynamics and $\gamma_{\mathsf{SS}(\sigma)}$ the spectral gap of the systematic scan with permutation $\sigma$. Corollary 3.1 instantiated with the Gibbs transition matrices implies
@@
The first identity implies
$\mathbb{E}\left[\sum_{\ell=j}^{n}k_{j}\right]$ $=\sum_{j=1}^{n}\sum_{\ell=1}^{j}\mathbb{E}[k_{\ell}-k_{\ell-1}]$
$=\sum_{j=1}^{n}\sum_{\ell=1}^{j}\frac{1}{p_{\ell}}$
$=\sum_{\ell=1}^{n}\frac{n-\ell+1}{p_{\ell}}$
$=n^{2}.$
These increments are also independent, and so
@@
We use similar notation as in Lemma 2.1. Fix any $f\in\mathcal{H}$ satisfying $\|f\|=1$ and define the sequence $0:=\gamma_{0}\leq\gamma_{1}\leq\ldots\leq\gamma_{L}$ via
@@
$\leq\sum_{\ell=1}^{k_{j}-1}\left\|\mathsf{Q}_{{\ell-1}}f-\mathsf{Q}_{{\ell}}f\right\|+\sum_{\ell=1}^{k_{j}}\left\|\mathsf{Q}_{{\ell}}f-\mathsf{Q}_{{\ell-1}}f\right\|$
@@
Taking squares, applying Cauchy-Schwarz, and telescoping via 8, we find that
@@
To see the last equality, note that
$\langle f,\mathsf{P}_{j}f-f\rangle=\langle f-\mathsf{P}_{j}f,\mathsf{P}_{j}f-f\rangle+\langle\mathsf{P}_{j}f,\mathsf{P}_{j}f-f\rangle=-\|\mathsf{P}_{j}f-f\|^{2},$
since any orthogonal projection is orthogonal to the remainder by the Pythagorean property. Finally, by assumption the left hand side is at most $(1-\delta)^{2}\leq 1-\delta$ so rearranging yields
@@ 3.2 From Scans to Glauber
We now show how a variation of the argument of Theorem 3.1 can be used to show a converse: if all, or even any, systematic scans have a spectral gap, then the Glauber dynamics has a spectral gap that is polynomially related. This provides a simple resolution to Open Problem 3.7 of Chlebicka, Łatuszyński, and Miasodejow [CŁM25]. However, unlike Theorem 1.1, we expect that our quantitative dependence is quantitatively suboptimal.
@@
$\mathsf{Q}_{{L-1}}f=\mathsf{Q}_{{k_{n-1}}}f+\sum_{\ell=k_{n-1}}^{L-2}(\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f),$
so a similar set of manipulations with the triangle inequality, Cauchy-Schwarz, and the fact projections can only decrease distance (noting $i_{k_{n-1}}=n-1$ by definition), and 11 yields
$\|\mathsf{Q}_{{L-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|$ $=\left\|\mathsf{Q}_{{k_{n-1}}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f+\sum_{\ell=k_{n-1}}^{L-2}(\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f)\right\|$
$\leq\|\mathsf{Q}_{{k_{n-1}}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|+\sum_{\ell=k_{n-1}}^{L-2}\|\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f\|$
$=\|\mathsf{P}_{n-1}\mathsf{Q}_{{k_{n-1}-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|+\sqrt{k_{n}-k_{n-1}-1}\sqrt{\sum_{\ell=k_{n-1}}^{L-2}\|\mathsf{Q}_{{\ell+1}}f-\mathsf{Q}_{{\ell}}f\|^{2}}$
$\leq\|\mathsf{Q}_{{k_{n-1}-1}}f-\mathsf{P}_{n-2}\ldots\mathsf{P}_{1}f\|+\sqrt{k_{n}-k_{n-1}-1}\sqrt{\sum_{\ell=k_{n-1}}^{L-2}(\lambda_{\ell+1}-\lambda_{\ell})}.$
@@
$\|\mathsf{Q}_{{L-1}}f-\mathsf{P}_{n-1}\ldots\mathsf{P}_{1}f\|$ $\leq\sum_{j=1}^{n-1}\sqrt{k_{j+1}-k_{j}-1}\sqrt{\sum_{\ell=k_{j}}^{k_{j+1}-2}(\lambda_{\ell+1}-\lambda_{\ell})}$
$\leq\sqrt{\sum_{j=1}^{n-1}(k_{j+1}-k_{j}-1)}\sqrt{\sum_{\ell\in S}(\lambda_{\ell+1}-\lambda_{\ell})}$
@@
where $S$ is a suitable subset of $[L]$; note that we use the fact that the inner summands for the increments are disjoint across different $k$ by construction. Putting this back with 13, we have
@@
$t^{\mathsf{GD}}_{\mathsf{mix}}(\varepsilon)\leq\max_{\sigma}\widetilde{O}\left(n^{4}(t_{\mathsf{mix}}^{\mathsf{SS}(\sigma)}(1/10))^{4}\right)\cdot\log\left(\frac{1}{\varepsilon\pi_{\min}}\right)$
$t^{\mathsf{GD}}_{\mathsf{mix}}(\varepsilon)\leq\min_{\sigma}\widetilde{O}\left(n^{6}(t_{\mathsf{mix}}^{\mathsf{SS}(\sigma)}(1/10))^{4}\right)\cdot\log\left(\frac{1}{\varepsilon\pi_{\min}}\right).$
@@ 4 Optimality of Results
Let $\pi$ be the uniform distribution on the set of spin configurations $\bm{x}\in\{0,1\}^{n}$ such that $\sum_{i=1}^{n}x_{i}\leq 1$. This is the so-called hardcore model of independent sets on the complete graph $K_{n}$ on $n$ nodes with fugacity $\lambda=1$. In particular, the set of configurations with positive (and uniform) probability is $\{\bm{0},\bm{e}_{1},\ldots,\bm{e}_{n}\}$, where $\bm{e}_{j}$ is the $j$th standard basis vector in $\mathbb{R}^{n}$. The associated Gibbs sampler is easily seen to act on configurations as follows when re-randomizing a spin $i$ at time $t$: if $X^{t}\in\{\bm{0},\bm{e}_{i}\}$, $X^{t+1}$ is uniform in $\{\bm{0},\bm{e}_{i}\}$, while otherwise $X^{t+1}=X^{t}$.
@@
If 16 holds, this implies the total variation between $c^{\prime}n^{2}$ full systematic scans and $\pi$ exceeds $1/2$ since $\pi(\{\bm{e}_{i}:i\in A\}\cup\{\bm{0}\})\leq 1/4$ while the dynamics has probability at least $3/4$ of being in this set.
To prove 16, observe that by independence of the geometric random variables in 15,
@@
$\Pr\left(\max_{s\in[s_{-},s_{+}]}\left|\sum_{i=1}^{s}Y_{i}-2s\right|\geq\lambda^{*}\right)\leq 1/10.$ (18)
Putting together 17 and 18 implies that
@@
Since $[2s_{-}-\lambda^{*},2s_{+}+\lambda^{*}]$ is a contiguous interval of length $O(\sqrt{c^{\prime}}n)$, the image $A$ of this set under the $\ (\mathrm{mod}\ n)$ map also has at most $n/4-1$ elements if $c^{\prime}$ is a small enough constant, and thus 19 demonstrates $A$ satisfies 16. By the discussion above, this verifies $t^{\mathsf{SS}}_{\mathrm{mix}}(1/4)\geq c^{\prime}n^{2}\geq cnt_{\mathrm{mix}}^{\mathsf{GD}}(1/4)$ so long as $c<c^{\prime}/C_{\mathsf{GD}}$.